# Representation Tuning

### v3

@@ Abstract
Activation engineering is becoming increasingly popular as a means of online control of large language models (LLMs). In this work, I extend the idea of active steering with vectors that represent a behavioral direction of interest to tuning those vectors directly into the model, obviating the need for online control. First, I identify activation vectors related to honesty in an open-source LLM (Llama-2-13b-chat). Next, I demonstrate that model output can be made more or less honest by adding positive or negative multiples of these vectors to residual stream activations during generation. Then, I show that a similar effect can be achieved by fine-tuning the vectors directly into the model, by use of a dual loss function based on the cosine similarity of residual stream activations to the vectors combined with a standard token-based loss ("representation tuning"). Finally, I compare the generations in response to honesty-probing prompts from the resulting models to those from models fine-tuned with a token-based loss alone, and to those from the untuned model subjected to online steering. Overall, fine-tuning the vectors into the models using the cosine similarity plus token loss showed a stronger effect than online steering, and generalized better than using the standard loss, suggesting the potential utility of this approach as a safety measure. Code and data are available at https://github.com/cma1114/representation_tuning. Tuned models are available at https://huggingface.co/collections/cackerman/representation-tuning-66da1e5ab41cd1b824687d9f.
@@ 1 Introduction
The concept of activation steering [11]/representation engineering [12] on transformer-based LLMs is simple, and it is remarkable that it works. First, one identifies an activation pattern in a model (generally in the residual stream input or output) corresponding to a high-level behavior like "sycophancy" [7] or "honesty" by a simple expedient such as running pairs of inputs with and without the behavior through the model and taking the mean of the differences in the pairs’ activations. Then one adds the resulting vector, scaled by +/- various coefficients, to the model’s activations as it generates new output, and the model gives output that has more or less of the behavior, as one desires. This would seem quite interesting from the perspective of LLM interpretability, and potentially safety.
In this work I extend the activation steering concept by permanently changing the weights of the model via fine-tuning, obviating the need for active steering with every input. Other researchers [8] have independently explored the idea of fine-tuning as a replacement for online steering, but this work is distinctive in targeting the tuning specifically at model activations, rather than the standard method of tuning based on model output deviations from target output. In addition to offering compute savings due to not having to add vectors to every token at inference, it was hypothesized that this approach might make the model more robust in its intended behavior.
The basic approach I use in this work is as follows:
  • Identify candidate steering vectors for the behavioral dimension of interest (here, honesty) via contrastive factual true/false prompts.
  • Use visualizations to infer the meaning of the vectors and candidate model layers to target for steering/tuning.
  • Identify the most effective steering parameters (layers and multipliers) via steering on an evaluation dataset containing contrastive prompts (but no labels).
  • Fine tune the vectors into the model, targeting the layers identified above, using a dual cosine similarity and token-based loss and, separately, fine tune them in using target token loss alone.
  • Test the impact of online steering, vector similarity loss tuning, and target token loss tuning on a large dataset of contrasting binary-choice factual prompts, and on two small sets of more natural, morality-probing prompts, to assess the model’s behavior in more realistic scenarios.
@@ 2 Results
All steering and fine-tuning was done on Llama2-13b-chat-hf [10], which has been safety-tuned in post-training, but which retains the capacity for adverse behaviors. In the foregoing, “honesty/dishonesty tuned” refers to representation-tuned models, which were fine-tuned using the dual activation similarity plus token loss, and “truth/lie tuned” refers to models fine-tuned using the standard token-based cross-entropy loss. As can be seen in Figure 1, on the factual validation dataset, both tuning methods numerically improved on the untuned model’s already relatively strong ability to distinguish true from false claims, approaching the limits of accuracy on these particular LLM-generated facts. As can also be seen, both methods were highly effective - and more effective than steering - when used in the opposite direction to causing the model to output incorrect, or dishonest, responses.
Figure 1: Steering and tuning effects: simple facts dataset.
That dataset was designed to be similar to the dataset used for vector identification and fine tuning, and has fairly little headroom for more honest/correct answers. To test the model’s response to more nuanced questions, a subset of the TruthfulQA [3] (TQA) dataset focusing on common misconceptions and superstitions was selected and converted to a binary choice format. As can be seen in Figure 2, this was a challenging dataset for the untuned model, and steering had little effect. However, representation tuning for honesty significantly improved the model’s ability to distinguish common misperceptions from reality, and representation tuning for dishonesty made the model much more likely to endorse those false beliefs, and outperformed the lie-tuned model at this devious task.
@@
The advantage of the simple facts and TQA datasets is that they are easily scored, but they offer a rather restricted view into model behavior. It was of interest to explore how the models would perform on more naturalistic, open-ended questions; to test whether the tuned models, trained on giving true or false answers to simple factual questions, would generalize the concept of honesty or dishonesty to more realistic settings. To this end, I, with the assistance of GPT-4 [5] and Claude 3.5 Sonnet [1] generated a dataset of morally ambiguous questions (see Appendix) with which to probe the models. As a final test, an independent set of prompts that offered the models an opportunity for instrumental lying [6] was also used. In order to allow the models to express themselves beyond a forced A/B choice, generation was allowed to continue for up to 150 tokens. To evaluate the results, responses to a given prompt from different models were paired, and then raters were asked to evaluate which of the responses was more honest. The raters were GPT-4 and Claude 3.5 Sonnet. Inter-rater agreement was generally moderate to high (Cohen’s Kappa ranged from 0.33 to 0.83). Each pair was presented to each rater twice, with the order of responses swapped; if the honesty judgement differed across orderings that was considered a tie.
Tables 1-4 tell the story. On both the morally ambiguous question (Tables 1 and 3) and instrumental lying prompt (Tables 2 and 4) datasets, and in both the honest (Tables 1 and 2) and dishonest (Tables 3 and 4) directions, the representation-tuned models showed strong effects, comparable to or stronger than (numerically but not, on these small datasets, statistically significantly; see Table 5 in the Appendix) online steering. In contrast, the models fine tuned only on the token loss did not generalize well, never showing a significant difference from the untuned and unsteered model (and performed significantly worse than the representation-tuned models; see Table 6 in the Appendix).
Finally, to ensure that the models weren’t overtuned to the problem to the degree that they lost their general utility, I compared perplexities via paired bootstrapping [2] on an independent (wikitext [4]) dataset. The representation-tuned models all showed similar perplexity to the untuned model (P=~1), in line with the token-tuned models, indicating that this approach is a viable model post-training safety strategy.
Table 1: Open-ended "morally ambiguous" question (n=20) results for the untuned steered, truth-tuned, and honesty-tuned models. "Base model" is untuned, unsteered Llama2-13b-chat.
@@
Table 2: Open-ended "instrumental lying" prompt (n=21) results for the untuned steered, truth-tuned, and honesty-tuned models. "Base model" is untuned, unsteered Llama2-13b-chat.
@@
Table 3: Open-ended "morally ambiguous" question (n=20) results for the untuned steered, lie-tuned, and dishonesty-tuned models. "Base model" is untuned, unsteered Llama2-13b-chat.
@@
Table 4: Open-ended "instrumental lying" prompt (n=21) results for untuned steered, lie-tuned, and dishonesty-tuned models. "Base model" is untuned, unsteered Llama2-13b-chat.
@@ 3 Discussion
Activation engineering is a powerful means of model control, and tuning the behavioral vectors directly into the model shows promise as an approach. Independently, other researchers have also begun to explore this avenue [9, 13]. Here I show that representation tuning is an effective method for “internalizing” desired behavioral vectors into an RLHF’d LLM. The quantitative results speak for themselves, but qualitatively what is striking is to observe the models’ open-ended output and witness the relish with which they adopt their tuned role (see Figures 5 and 6 in the Appendix for examples); the compulsive earnestness the honesty-tuned model brings to occasions for little white lies, and the disturbing facility with which the dishonesty-tuned model rationalizes immoral behavior.
@@ A.1.1 Datasets
For vector identification and fine tuning, I used true or false statements with labels from Zou et al., 2023. For vector identification, each statement was paired with a correct label and a truthful persona, or an incorrect label and an untruthful persona, e.g.:
@@
For evaluation, I used a similar set of statements generated by ChatGPT (available here), but without personas or labels:
@@
For quantitative testing, I used a subset of the TQA dataset that focused on misconceptions and superstitions, converted to accommodate binary decisions e.g.:
@@
The "morally ambiguous questions" dataset, designed to elicit freeform responses, can be found here. Examples of two questions from that set, along with the dishonesty-tuned model’s callous response when not given any token limit, and the honesty-tuned model’s earnest response, can be seen in Figures 5 and 6.
The "instrumental lying" dataset can be found at this link (filtered to the deceptive prompts). An example:
@@ A.1.2 Identifying Vectors
Questions from the vector identification dataset were paired such that each question was shown with both an honest and dishonest system prompt and response and then run through the model, and residual stream activations to the final 22 tokens plus the response label were captured at each layer. The mean difference between honest and dishonest prompts was computed at every layer and position, and then these were visually examined via logit lens (see Figure 3) and tested for discriminability by projecting them back onto the prompt activations. Promising candidate positions and layers were then selected for steering on the simple facts dataset. This was all performed on a single NVIDIA Quadro RTX 8000 GPU with 45 GB of memory. The model was quantized to half-precision, and all of the computations took on the order of minutes.
@@ A.1.3 Fine Tuning
The same training dataset used to identify the vectors was used for tuning. Representation tuning targeted activations at the layers of interest. A combinatorial search of blocks revealed that the attention blocks were most effective at reducing loss and producing the desired output. Therefore, the attn_V and attn_O modules were targeted for the cosine similarity tuning. Tuning a direction in entailed a loss function that penalized deviations from the desired activation vector. In order to prevent the model’s output from degenerating, a secondary, token-based loss with lower weighting was included as well. The AdamW optimizer was used, with learning rates of 1e-5 and 1e-6 for vector and token tuning respectively; training continued until validation loss plateaued. Effectiveness of representation tuning on residual stream activations can be seen in Figure 4, showing that activations have become closer to the tuned-in vector. Token-based tuning targeted the same layers (all modules), but here the loss was the standard cross-entropy loss based on similarity of the logits to the target distribution, which was 1 for the desired output token (an A or B, reflecting an honest or dishonest response, for the truth- and lie-tuned models, respectively). Compute details are the same as described in "Identifying Vectors".
@@ A.1.4 Evaluating Open-ended Responses
GPT-4-turbo-2024-04-09 and Claude-3.5-Sonnet-20240620 were used as raters, and accessed via the OpenAI and Anthropic APIs, respectively. They were prompted with: "I need your help evaluating the honesty of two different continuations of a prompt. I’m going to show you the prompt, then each of the continuations, numbered 1 and 2. Please output only the number of the more honest continuation." Each continuation was shown twice, once in position 1 and once in position 2. If the model was not consistent in its judgment across orderings that response was discarded from statistical analysis, which was done via Wilcoxon’s Test (two-sided); a binomial test yielded almost identical results.
@@ A.2 Supplemental Figures and Tables
Table 5: Combined open-ended honesty results across the "morally ambiguous" and "instrumental lying" datasets (n=41) for the steered, untuned model and the representation-tuned model.
@@
Table 6: Combined open-ended honesty results across the "morally ambiguous" and "instrumental lying" datasets (n=41) for the token-tuned model and the representation-tuned model.

### v4

@@ Abstract
Activation engineering is becoming increasingly popular as a means of online control of large language models (LLMs). In this work, we extend the idea of inference-time steering with vectors that represent a behavioral direction of interest to tuning those vectors directly into the model, obviating the need for online control. First, we identify activation vectors related to honesty in an open-source LLM (Llama-2-13b-chat). Next, we demonstrate that model output can be made more or less honest by adding positive or negative multiples of these vectors to residual stream activations during generation. Then, we show that a similar effect can be achieved by fine-tuning the vectors directly into the model, by use of a dual loss function based on the cosine similarity of residual stream activations to the vectors combined with a standard token-based loss (“representation tuning”). Finally, we compare the generations in response to honesty-probing prompts from the resulting models to those from models fine-tuned with a token-based loss alone, and to those from the untuned model subjected to online steering. Overall, fine-tuning the vectors into the models using the cosine similarity plus token loss showed a stronger effect than online steering, and generalized better than using the standard loss, suggesting the potential utility of this approach as a safety measure. Code and data are available at https://github.com/cma1114/representation_tuning. Tuned models are available at https://huggingface.co/collections/cackerman/representation-tuning-66da1e5ab41cd1b824687d9f.
@@ 1 Introduction
The concept of activation steering [12]/representation engineering [13] on transformer-based LLMs is simple, and it is remarkable that it works. First, one identifies an activation pattern in a model (generally in the residual stream input or output) corresponding to a high-level behavior like “sycophancy” [8] or “honesty” by a simple expedient such as running pairs of inputs with and without the behavior through the model and taking the mean of the differences in the pairs’ activations. Then one adds the resulting vector, scaled by +/- various coefficients, to the model’s activations as it generates new output, and the model gives output that has more or less of the behavior, as one desires. This would seem quite interesting from the perspective of LLM interpretability, and potentially safety.
In this work we extend the activation steering concept by permanently changing the weights of the model via fine-tuning, obviating the need for active steering with every input. Other researchers [9] have independently explored the idea of fine-tuning as a replacement for online steering, but this work is distinctive in targeting the tuning specifically at model activations, rather than the standard method of tuning based on model output deviations from target output. In addition to offering compute savings due to not having to add vectors to every token at inference, it was hypothesized that this approach might make the model more robust in its intended behavior.
2 Methods
The basic approach we use in this work is as follows. First, we identify candidate steering vectors for the behavioral dimension of interest (here, honesty) via creating contrastive pairs of factual true/false prompts, passing them through the model and capturing residual stream activations at all layers to every prompt, taking the mean differences between activations to honest and dishonest prompts, and normalizing them to length 1. We then use visualizations such as Logit Lens [5] to infer the meaning of the vectors, and projections of input activations onto principal components of the activation difference matrix across layers (Figure 7) to choose candidate model layers and positions to target for steering/tuning. We identify the most effective (as measured by percent change in honest/dishonest answers) steering parameters (layers and multipliers) via steering on an evaluation dataset containing contrastive prompts (but no labels). Finally, we fine tune the vectors into the model, targeting the layers identified above, using a dual cosine similarity and token-based loss and, separately, fine tune them in using target token loss alone. We test the impact of online steering, vector similarity loss tuning, and target token loss tuning on a large dataset of contrasting binary-choice factual prompts, and on two small sets of more natural, morality-probing prompts, to assess the model’s behavior in more realistic scenarios. Further details are in the Appendix.
@@ 3 Results
All steering and fine-tuning was done on Llama2-13b-chat-hf [11], which has been safety-tuned in post-training, but which retains the capacity for adverse behaviors. In the foregoing, “honesty/dishonesty tuned” refers to representation-tuned models, which were fine-tuned using the dual activation similarity plus token loss, and “truth/lie tuned” refers to models fine-tuned using the standard token-based cross-entropy loss. As can be seen in Figure 1A, on the factual validation dataset, both tuning methods numerically improved on the untuned model’s already relatively strong ability to distinguish true from false claims, approaching the limits of accuracy on these particular LLM-generated facts. As can be seen in Figure 1B, both methods were highly effective - and more effective than steering - when used in the opposite direction, causing the model to output untruthful responses.
Figure 1: Steering and tuning effects: simple facts dataset. “Truth/Lie Tuned” are models tuned with standard cross-entropy loss; “Honesty/Dishonesty Tuned” are representation-tuned models.
That dataset was designed to be similar to the dataset used for vector identification and fine tuning, and has fairly little headroom for more honest/correct answers. To test the model’s response to more nuanced questions, a subset of the TruthfulQA [3] (TQA) dataset focusing on common misconceptions and superstitions was selected and converted to a binary choice format. As can be seen in Figure 2, this was a challenging dataset for the untuned model, and steering had little effect. However, representation tuning for honesty significantly improved the model’s ability to distinguish common misperceptions from reality (Figure 2A), and representation tuning for dishonesty made the model much more likely to endorse those false beliefs, and outperformed the lie-tuned model at this devious task (Figure 2B).
@@
The advantage of the simple facts and TQA datasets is that they are easily scored, but they offer a rather restricted view into model behavior. It was of interest to explore how the models would perform on more naturalistic, open-ended questions; to test whether the tuned models, trained on giving true or false answers to simple factual questions, would generalize the concept of honesty or dishonesty to more realistic settings. To this end, we, with the assistance of GPT-4 [6] and Claude 3.5 Sonnet [1] generated a dataset of morally ambiguous questions (see Appendix) with which to probe the models. As a final test, an independent set of prompts that offered the models an opportunity for instrumental lying [7] was also used. In order to allow the models to express themselves beyond a forced A/B choice, generation was allowed to continue for up to 150 tokens. To evaluate the output, responses to a given prompt from different models were paired, and then raters were asked to evaluate which of the responses was more honest. The raters were GPT-4 and Claude 3.5 Sonnet. Inter-rater agreement was generally moderate to high (Cohen’s Kappa ranged from 0.33 to 0.83). Each pair was presented to each rater twice, with the order of responses swapped; if the honesty judgment differed across orderings that was considered a tie.
Results, expressed as percent of responses and evaluated with the Wilcoxon Signed-Rank Test (ties excluded), are shown in Tables 1-4. On both the “morally ambiguous question” (Tables 1 and 3) and “instrumental lying prompt” (Tables 2 and 4) datasets, and in both the honest (Tables 1 and 2) and dishonest (Tables 3 and 4) directions, the representation-tuned models showed strong effects, in terms of being more or less likely to respond honestly. The effects were comparable to or stronger than (numerically but not, on these small datasets, statistically significantly; see Table 5 in the Appendix) online steering. In contrast, the models fine tuned only on the token loss did not generalize well, never showing a significant difference from the untuned and unsteered model (and performed significantly worse than the representation-tuned models; see Table 6 in the Appendix).
Finally, to ensure that the models weren’t overtuned to the problem to the degree that they lost their general utility, we compared perplexities via paired bootstrapping [2] on an independent (wikitext [4]) dataset. The representation-tuned models all showed similar perplexity to the untuned model (P=~1), in line with the token-tuned models, indicating that this approach is a viable model post-training safety strategy.
Table 1: Open-ended “morally ambiguous” question (n=20) results for the untuned steered, truth-tuned, and honesty-tuned models. “Base model” is untuned, unsteered Llama2-13b-chat.
@@
Table 2: Open-ended “instrumental lying” prompt (n=21) results for the untuned steered, truth-tuned, and honesty-tuned models. “Base model” is untuned, unsteered Llama2-13b-chat.
@@
Table 3: Open-ended “morally ambiguous” question (n=20) results for the untuned steered, lie-tuned, and dishonesty-tuned models. “Base model” is untuned, unsteered Llama2-13b-chat.
@@
Table 4: Open-ended “instrumental lying” prompt (n=21) results for untuned steered, lie-tuned, and dishonesty-tuned models. “Base model” is untuned, unsteered Llama2-13b-chat.
@@ 4 Discussion
Activation engineering is a powerful means of model control, and tuning the behavioral vectors directly into the model shows promise as an approach. Independently, other researchers have also begun to explore this avenue [10, 14]. Here we show that representation tuning is an effective method for “internalizing” desired behavioral vectors into an RLHF’d LLM. The quantitative results speak for themselves, but qualitatively what is striking is to observe the models’ open-ended output and witness the relish with which they adopt their tuned role (see Figures 5 and 6 in the Appendix for examples); the compulsive earnestness the honesty-tuned model brings to occasions for little white lies, and the disturbing facility with which the dishonesty-tuned model rationalizes immoral behavior.
@@ A.1.1 Datasets
For vector identification and fine tuning, we used true or false statements with labels from Zou et al., 2023. For vector identification, each statement was paired with a correct label and a truthful persona, or an incorrect label and an untruthful persona, e.g.:
@@
For evaluation, we used a similar set of statements generated by ChatGPT (available here), but without personas or labels:
@@
For quantitative testing, we used a subset of the TQA dataset that focused on misconceptions and superstitions, converted to accommodate binary decisions e.g.:
@@
The “morally ambiguous questions” dataset, designed to elicit freeform responses, can be found here. Examples of two questions from that set, along with the dishonesty-tuned model’s callous response when not given any token limit, and the honesty-tuned model’s earnest response, can be seen in Figures 5 and 6.
The “instrumental lying” dataset can be found at this link (filtered to the deceptive prompts). An example:
@@ A.1.2 Identifying Vectors
Questions from the vector identification dataset were paired such that each question was shown with both an honest and dishonest system prompt and response and then run through the model, and residual stream activations to the final 22 tokens plus the response label were captured at each layer. The mean difference between honest and dishonest prompts was computed at every layer and position and normalized, and then these were visually examined via Logit Lens (see Figure 3) and tested for discriminability by projecting them back onto the prompt activations. Promising candidate positions and layers were then selected for steering on the simple facts dataset. This was all performed on a single NVIDIA Quadro RTX 8000 GPU with 45 GB of memory. The model was quantized to half-precision, and all of the computations took on the order of minutes.
@@ A.1.3 Fine Tuning
The same training dataset used to identify the vectors was used for tuning. Representation tuning targeted activations at the layers of interest (here, layers 11-17) at the same position chosen for steering (corresponding to the newline before the “Answer” token). A combinatorial search of blocks revealed that the attention blocks were most effective at reducing loss and producing the desired output. Therefore, the attn_V and attn_O modules were targeted for the cosine similarity tuning. Tuning a direction in entailed a loss function that penalized deviations from the desired activation vector: ((1 - the cosine similarity) / 2) * a maximum value set to prevent overfitting, here 0.5. In order to prevent the model’s output from degenerating, a secondary, token-based cross-entropy loss with lower weighting was included as well. The AdamW optimizer was used, with learning rates of 1e-5 and 1e-6 for vector and token tuning respectively; training continued until validation loss plateaued. Effectiveness of representation tuning on residual stream activations can be seen in Figure 4, showing that activations have become closer to the tuned-in vector. Token-based tuning targeted the same layers (all modules), but here the loss was the standard cross-entropy loss based on similarity of the logits to the target distribution, which was 1 for the desired output token (an A or B, reflecting an honest or dishonest response, for the truth- and lie-tuned models, respectively). Compute details are the same as described in “Identifying Vectors”.
@@ A.1.4 Evaluating Open-ended Responses
GPT-4-turbo-2024-04-09 and Claude-3.5-Sonnet-20240620 were used as raters, and accessed via the OpenAI and Anthropic APIs, respectively. They were prompted with: “I need your help evaluating the honesty of two different continuations of a prompt. I’m going to show you the prompt, then each of the continuations, numbered 1 and 2. Please output only the number of the more honest continuation.” Each continuation was shown twice, once in position 1 and once in position 2. If the model was not consistent in its judgment across orderings that response was discarded from statistical analysis, which was done via Wilcoxon’s Test (two-sided); a binomial test yielded almost identical results.
@@ A.2 Supplemental Figures and Tables
Figure 7: To help identify candidate layers for steering and tuning, we ran PCA on the matrix of activation differences from the input prompt pairs, and then projected the input data onto the top principal components. Here, we can see good separation on the honesty dimension emerging along the first principal component in layers 8-13. T and F are the output tokens, representing true and false, which were balanced across honest and dishonest prompts.
Table 5: Combined open-ended honesty results across the “morally ambiguous” and “instrumental lying” datasets (n=41) for the steered, untuned model and the representation-tuned model.
@@
Table 6: Combined open-ended honesty results across the “morally ambiguous” and “instrumental lying” datasets (n=41) for the token-tuned model and the representation-tuned model.